# Solving the QCD effective kinetic theory with neural networks

### v1

@@ I Introduction
Crucially, while the collision kernel is non-local in the momentum space, it remains local in coordinate space. Then, at any given space-time point, the collision kernel is a deterministic, non-linear map from the distribution function to its time derivative, which can be, in principle, fitted or parametrized with a sufficiently versatile fitting function. Here, we develop such a fit by training a neural network with a large amount of training data obtained by using the traditional Monte Carlo evaluation of the collision kernel Xiao and Frank (2021, 2023). The benefit of this approach is that once the neural-network model of the collision kernel is trained, it can be used very efficiently in any time evolution of complicated event-by-event time-evolutions instead of spending large amounts of computing resources on repeated evaluations of distribution functions that closely resemble each other, both across different space-time points and across events. In fact, it has been observed that the phenomenon of non-equilibrium attractors Heller and Spalinski (2015); Kurkela et al. (2020b); Almaalol et al. (2020); Boguslavski et al. (2024) reduces the relevant space of relevant distribution functions, greatly simplifying the functional space needed to cover when training the neural network.
@@ II.1 Collision kernel in EKT
Here, the velocity is $v^{\mu}=p^{\mu}/p^{0}$. The left-hand side accounts for the free-streaming space-time evolution and is local in $\mathrm{\mathbf{p}}$-space. The collision kernel $\mathcal{C}[f(x^{\mu},\mathrm{\mathbf{p}})]$ of the right-hand side accounts for the scattering between the partons, and is local in $\mathrm{\mathbf{x}}$-space. It consists of elastic $2\leftrightarrow 2$–scattering and medium-induced (effective) collinear $1\leftrightarrow 2$ splitting/merging processes
@@
×$\quad\times(2\pi)^{4}\delta^{4}(P+K-P^{\prime}-K^{\prime})$ (3)
×$~{}~{}~{}\times\Big{\{}f(\mathrm{\mathbf{p}})f(\mathrm{\mathbf{k}})\left[1+f(\mathrm{\mathbf{p}}^{\prime})\right]\left[1+f(\mathrm{\mathbf{k}}^{\prime})\right]$
$~{}~{}\quad-f(\mathrm{\mathbf{p}}^{\prime})f(\mathrm{\mathbf{k}}^{\prime})\left[1+f(\mathrm{\mathbf{p}})\right]\left[1\pm f(\mathrm{\mathbf{k}})\right]\Big{\}},$
where $\int_{\mathrm{\mathbf{k}}}=\int\frac{\mathrm{d}^{3}{\mathrm{\mathbf{k}}}\,}{(2\pi)^{3}2|\mathrm{\mathbf{k}}|}$ denotes the Lorentz-invarant integration measure for on-shell momenta ($K^{0}=|\mathrm{\mathbf{k}}|$), and $\nu=2d_{A}$ counts the number of degrees of freedom for gluons. For simplicity, we have suppressed the space-time coordinate $x^{\mu}$ in all distribution functions, since it is local. For the matrix element we use the leading-order pQCD $gg-gg$ scattering matrix element summed over all incoming and outgoing spin and color degrees of freedom. We include medium effects by screening the divergent $t$ and $u$ channel propagators using the isotropic HTL (isoHTL) propagators for the internal soft gluon lines Arnold et al. (2003a); Boguslavski and Lindenbauer (2024). We use the ’t Hooft coupling $\lambda=g^{2}N_{C}=10$ throughout this paper.
@@
[…]
With only two momenta and strictly collinear splitting, its kinematics is simpler than for the elastic collision term. However, the effective matrix elements for the splitting/merging processes $\gamma$ are more complicated since they need to account for the quantum-mechanical formation time of the splitting-merging process, and effectively interpolate between the Bethe-Heitler and LPM rates Arnold et al. (2002). We describe the collision terms in more detail in the Appendix A.
@@ II.2 Collision Kernel as a machine-learning problem
More generally, we can view the collision kernel as a map of the space of distribution functions to their time derivatives $\mathcal{C}:f\mapsto\partial_{t}f$. After discretizing momentum space $\mathrm{\mathbf{p}}=(p_{i},\cos\theta_{j},\phi_{k})$, we can represent the distribution function at every space-time point as a $n_{X}$-dimensional vector, $f(\mathrm{\mathbf{p}})\in\mathbb{R}^{n_{X}}$. The collision kernel is then a nonlinear map
@@
with $n_{X}$ being the number of points in the discrete momentum space. Modeling such a map is a task well-suited to a perceptron network Hornik et al. (1989); Mehta et al. (2019).
The collision kernel can be computed using Monte Carlo integration techniques, as detailed in Appendix A. However, this is costly, and many of the distribution functions closely resemble each other. Therefore, in a fully 3+3+1D system (corresponding to a distribution function $f(\mathrm{\mathbf{x}},\mathrm{\mathbf{p}},t)$), we need to repeatedly obtain the collision kernel for very similar distribution functions.
@@ II.3 Conformal symmetry of the EKT
In our implementation of the collision kernel, we work with dimensionless quantities where the temperature is scaled out. The dimensionless quantities are denoted with a hat, i.e.,
$\hat{p}=p/T,$ $\hat{t}=tT,$ $\hat{T}=1,$ $\hat{\varepsilon}=\pi^{2}/30.$ (8)
@@ II.4 Discretization
and similarly for the polar ($\cos\theta$) and azimuthal ($\phi$) grid. At any given momentum $\mathrm{\mathbf{p}}$, the distribution function $f(\mathrm{\mathbf{p}})$ is encoded in the neighbouring grid points $(p_{i},\cos\theta_{j},\phi_{k})$. At the grid points themselves, the relation is approximately
@@
Between the grid points, $f(\mathrm{\mathbf{p}})$ is found by linearly interpolating the neighboring results. Important discretization parameters are the grid boundaries for the $p$-grid, $p_{\mathrm{min}}$, and $p_{\mathrm{max}}$.
@@ II.5 Data preprocessing and the coarse grid
To generate training data using the Monte Carlo method, a fine grid with a hundred or more points per dimension must be employed, as the method of Kurkela and Moore (2012) suffers from numerical instabilities during time evolution—instabilities that are mitigated by finer grid spacing. These instabilities can arise when the collision kernel is evaluated between grid points where one is overoccupied and the other underoccupied. However, despite the fine grid needed for the Monte Carlo method, the physical information can often be reduced to a small number of parameters, e.g., peak positions, width, and weight; or, more physically, moments of the distribution functions, which are used, for instance, in hydrodynamical models Denicol et al. (2012). Hence, some of the information stored in the fine grid in the Monte Carlo method is redundant, and for a neural network it may be sufficient to use a dimensionally reduced input data set. In particular, for the reasonable accuracy needed for phenomenological applications, a very fine grid is also not needed.
@@ III.1 Generation of training data, choice of initial conditions
$g(\hat{p})=\begin{cases}1&\text{almost everywhere}\\ <1&\text{for a very small region}\end{cases}$ (18)
@@
$g(\hat{p})=1-a_{2}e^{-s_{2}(\hat{p}-\hat{p}_{2})^{2}},$ (19)
with $a_{2}$, $p_{2}$, and $s_{2}$ chosen such that the total energy is conserved, with $10<a_{2}<$ 1.
@@
A representative sample of the near- and far-from-equilibrium distribution functions and their corresponding collision kernels in the training data set is displayed in Fig. 2.
@@ III.2 Hyperparameter optimization and uncertainty quantification
Many different neural network architectures have been described and proposed in the literature, which are designed for different tasks and purposes. Here, we use a feed-forward neural network with linear layers to obtain the collision kernel. For the precise architecture, i.e., number of layers and nodes per layer, we use an automated optimization framework, Ray Tune Liaw et al. (2018), to find the optimal set of hyperparameters. This hyperparameter tuning has the freedom to explore models with up to 5 layers; each of these can have 16, 32, 64, 128, 192, 256, 384, or 512 neurons.
We allow Ray Tune to choose not only the architecture of our neural network but also the learning rate as an additional hyperparameter. The architectures and learning rates with the best performance for each collision kernel are presented in Table 1. This tuning is performed for both collision kernels separately.
@@
Table 1: Best hyperparameters found for each collision kernels. The activation function of the neural network is the ReLU function in all layers.
@@ III.3 Results
The training of the 20 neural networks that we performed for this case involved around 200000 samples, which translates to a volume of data of $\sim 100~{}\mathrm{MB}$ [^2] We provide the data in Ref. Barrera Cabodevila et al. (2025), which includes the raw data and preprocessing steps, effectively amounting to more than the $100~{}\mathrm{MB}$ of data. To evaluate the performance of the described approach, we compare the evolution of several initial conditions with the results obtained within the Monte Carlo approach. In Fig. 4 we present the distribution functions and their corresponding collision kernels at different times for three different evolutions. The first two are not contained in the training data, while the third one has been extracted from the training dataset. We observe excellent agreement between the two approaches. All the Monte Carlo results lie within the error bands obtained from the neural network evolution and Eq. (21).
@@ IV.4 Results
The first index $n\geq 0$ labels the power of $p$ in the integrand. The indices $l$ and $m$ are the indices belonging to the spherical harmonics $Y_{l}^{m}(\theta,\phi)$, and obey the condition $l\geq 0$ and $-l\leq m\leq l$. Note that the Debye mass is proportional to the moment $M_{000}$, the number density to $M_{100}$, the energy density to $M_{200}$. Other familiar quantities can also be obtained from these moments. For instance, the pressure $P_{z}$ is given by $P_{z}/T^{4}=\frac{1}{3}M_{200}+\frac{2}{3}M_{220}$.
@@
We provide additional comparisons for simulations with different initial conditions in Fig. 7. There, we again first note the good overall agreement between the neural network and the Monte Carlo evolution. However, especially in the right panel, the error bars significantly grow during the evolution. Especially close to equilibrium, the network seems to struggle to remain and approach the thermal fixed point. Thus, despite enlarging the training data with a large amount of static distributions close to equilibrium, the neural network fails to accurately learn the approach to thermal equilibrium. This indicates that even more training data might be needed, or a different network architecture that by construction conserves the thermal fixed point.
@@ Acknowledgements.
We would like to thank Andreas Ipp, Kayran Schmidt, and David I. Müller for insightful discussions. FL is supported by the Austrian Science Fund (FWF) under Grant DOI 10.55776/P34455. FL is a recipient of a DOC Fellowship of the Austrian Academy of Sciences at TU Wien (project 27203). The results in this paper have been achieved in part using the Vienna Scientific Cluster (VSC), project 71444. SBC is supported by the European Research Council project ERC-2018-ADG-835105 YoctoLHC; by María de Maeztu grant CEX2023-001318-M and by project PID2023-152762NB-I00, both funded by MCIN/AEI/10.13039/-501100011033; from the Xunta de Galicia (CIGUS Network of Research Centres) and the European Union. SBC also acknowledges the support of Axudas de apoio á etapa predoutoral program (Ref. ED481A 2022/279).

### v2

@@ I Introduction
Crucially, while the collision kernel is non-local in the momentum space, it remains local in coordinate space. Then, at any given space-time point, the collision kernel is a deterministic, non-linear map from the distribution function to its time derivative, which can be, in principle, fitted or parametrized with a sufficiently versatile fitting function. Here, we develop such a fit by training a neural network with a large amount of training data obtained by using the traditional Monte Carlo evaluation of the collision kernel Xiao and Frank (2021, 2023). The benefit of this approach is that once the neural-network model of the collision kernel is trained, it can be used very efficiently in any time evolution of complicated event-by-event time-evolutions instead of spending large amounts of computing resources on repeated evaluations of distribution functions that closely resemble each other, both across different space-time points and across events. In fact, it has been observed that the phenomenon of non-equilibrium attractors Heller and Spalinski (2015); Kurkela et al. (2020b); Almaalol et al. (2020); Boguslavski et al. (2024) reduces the relevant space of relevant distribution functions, greatly simplifying the functional space needed to cover when training the neural network when applied to expanding kinetic theory simulations.
@@ II.1 Collision kernel in EKT
Here, the velocity is $v^{\mu}=p^{\mu}/p^{0}$. The left-hand side accounts for the free-streaming space-time evolution and is local in $\mathrm{\mathbf{p}}$-space but is not local in $\mathrm{\mathbf{x}}$-space in the sense that it connects neighboring points by derivatives. The collision kernel $\mathcal{C}[f(x^{\mu},\mathrm{\mathbf{p}})]$ of the right-hand side accounts for the scattering between the partons, and is strictly local in $\mathrm{\mathbf{x}}$-space but nonlocal in $\mathrm{\mathbf{p}}$-space. It consists of elastic $2\leftrightarrow 2$–scattering and medium-induced (effective) collinear $1\leftrightarrow 2$ splitting/merging processes
@@
$\quad\times(2\pi)^{4}\delta^{4}(P+K-P^{\prime}-K^{\prime})$ (3)
$~~~\times\Big\{f(\mathrm{\mathbf{p}})f(\mathrm{\mathbf{k}})\left[1+f(\mathrm{\mathbf{p}}^{\prime})\right]\left[1+f(\mathrm{\mathbf{k}}^{\prime})\right]$
$~~\quad-f(\mathrm{\mathbf{p}}^{\prime})f(\mathrm{\mathbf{k}}^{\prime})\left[1+f(\mathrm{\mathbf{p}})\right]\left[1+f(\mathrm{\mathbf{k}})\right]\Big\},$
where $\int_{\mathrm{\mathbf{k}}}=\int\frac{\mathrm{d}^{3}{\mathrm{\mathbf{k}}}\,}{(2\pi)^{3}2|\mathrm{\mathbf{k}}|}$ denotes the Lorentz-invarant integration measure for on-shell momenta ($K^{0}=|\mathrm{\mathbf{k}}|$), and $\nu=2d_{A}{=2(N_{c}^{2}-1)}$ counts the number of degrees of freedom for gluons. For simplicity, we have suppressed the space-time coordinate $x^{\mu}$ in all distribution functions, since it is local. For the matrix element we use the leading-order pQCD gggggg-$g$ scattering matrix element summed over all incoming and outgoing spin and color degrees of freedom. We include medium effects by screening the divergent $t$ and $u$ channel propagators using the isotropic HTL (isoHTL) propagators for the internal soft gluon lines Arnold et al. (2003a); Boguslavski and Lindenbauer (2024). We use the ’t Hooft coupling $\lambda=g^{2}N_{C}=10$ throughout this paper.
@@
$\begin{split}\mathcal{C}^{1\leftrightarrow 2}[f(\mathrm{\mathbf{\tilde{p}}})]&=\frac{(2\pi)^{3}}{4\pi\tilde{p}^{2}}\frac{1}{\nu}\int_{0}^{\infty}\mathrm{d}{p}\,\int_{0}^{p/2}\mathrm{d}{k^{\prime}}\,4\pi\gamma^{p}_{p^{\prime},k^{\prime}}\\ &\quad\times\Big\{f(\mathrm{\mathbf{p}})(1+f(p^{\prime}\hat{\mathrm{\mathbf{p}}}))(1+f(k^{\prime}\hat{\mathrm{\mathbf{p}}}))\\ &\qquad\quad-f(p^{\prime}\hat{\mathrm{\mathbf{p}}})f(k^{\prime}\hat{\mathrm{\mathbf{p}}})(1+f(\mathrm{\mathbf{p}}))\Big\}\\ &\times\left[\delta(\tilde{p}-p)-\delta(\tilde{p}-p^{\prime})-\delta(\tilde{p}-k^{\prime})\right]\end{split}$ (4)
With only two momenta and strictly collinear splitting in the direction of $\mathrm{\mathbf{p}}$ (formalized by the unit vector $\hat{\mathrm{\mathbf{p}}}=\mathrm{\mathbf{p}}/p$), its kinematics is simpler than for the elastic collision term. However, the effective matrix elements for the splitting/merging processes $\gamma$ are more complicated since they need to account for the quantum-mechanical formation time of the splitting-merging process, and effectively interpolate between the Bethe-Heitler and LPM rates Arnold et al. (2002). We describe the collision terms in more detail in the Appendix A.
@@ II.2 Collision Kernel as a machine-learning problem
More generally, we can view the collision kernel as a map of the space of distribution functions to their time derivatives $\mathcal{C}:f\mapsto(\partial_{t}f)_{\mathrm{coll}}$. After discretizing momentum space $\mathrm{\mathbf{p}}=(p_{i},\cos\theta_{j},\phi_{k})$, we can represent the distribution function at every space-time point as a $n_{X}$-dimensional vector, $f(\mathrm{\mathbf{p}})\in\mathbb{R}^{n_{X}}$. The collision kernel is then a nonlinear map
@@
with $n_{X}$ being the number of points in the discrete momentum space. Modeling such a map is a task well-suited to an artificial neural network Hornik et al. (1989); Mehta et al. (2019).
The collision kernel can be computed using Monte Carlo integration techniques, as detailed in Appendix A. However, this is costly, and many of the distribution functions closely resemble each other. Therefore, in a fully 3+3+1D system (corresponding to a distribution function $f(\mathrm{\mathbf{x}},\mathrm{\mathbf{p}},t)$), we need to repeatedly obtain the collision kernel for very similar distribution functions. However, since the collision kernels are local in $\mathrm{\mathbf{x}}$-space, their calculation only involves the value of the distribution function in the current cell of the spatial grid. In this work, we demonstrate that we can compute these collision kernels efficiently with machine learning techniques. For this reason, we don’t include the spatial dependence in our results and we leave the full 3+3+1D simulation for the future.
@@ II.3 Conformal symmetry of the EKT
In our implementation of the collision kernel, we work with dimensionless quantities where the temperature is scaled out.
@@ II.4 Discretization
and the boundary bins are constructed as $w_{0}(p)=(p_{1}-p)/(p_{1}-p_{0})$ for $p_{0}<p<p_{1}$ and zero otherwise, and $w_{n}(p)=(p-p_{n-1})/(p_{n}-p_{n-1})$ for $p_{n-1}<p<p_{n}$ and zero otherwise. The grid for the polar ($\cos\theta$) and azimuthal ($\phi$) angles is constructed similarly, with the exception that the polar angle grid is periodic, and thus there are no boundary bins. At any given momentum $\mathrm{\mathbf{p}}$, the distribution function $f(\mathrm{\mathbf{p}})$ is encoded in the neighbouring grid points $(p_{i},\cos\theta_{j},\phi_{k})$. At the grid points themselves, the relation is approximately
@@
Between the grid points, $f(\mathrm{\mathbf{p}})$ is found by linearly interpolating the neighboring results. Important discretization parameters are the grid boundaries for the $p$-grid, $p_{\mathrm{min}}$, and $p_{\mathrm{max}}$. In the presented results, we always choose $p_{\mathrm{max}}=15T$. However, the $p_{\mathrm{min}}$ value is different for the isotropic and anisotropic cases. We set $p_{\mathrm{min}}=0.01T$ in the former and $p_{\mathrm{min}}=0.1T$ in the latter.
@@ II.5 Data preprocessing and the coarse grid
To generate training data using the Monte Carlo method, a fine grid with typically a large number of points per dimension must be employed, as the method of Abraao York et al. (2014) suffers from numerical instabilities during time evolution–—instabilities that are mitigated by finer grid spacing. These instabilities can arise when the collision kernel is evaluated between grid points where one is overoccupied and the other underoccupied. However, despite the fine grid needed for the Monte Carlo method, the physical information can often be reduced to a small number of parameters, e.g., peak positions, width, and weight; or, more physically, moments of the distribution functions, which are used, for instance, in hydrodynamical models Denicol et al. (2012). Hence, some of the information stored in the fine grid in the Monte Carlo method is redundant, and for a neural network it may be sufficient to use a dimensionally reduced input data set. In particular, for the reasonable accuracy needed for phenomenological applications, a very fine grid is also not needed.
@@ III.1 Generation of training data, choice of initial conditions
$g(p{/T})=\begin{cases}1&\text{almost everywhere}\\ <1&\text{for a very small region}\end{cases}$ (17)
@@
$g(p{/T})=1-a_{2}e^{-s_{2}(p-p_{2})^{2}{/T^{2}}},$ (18)
with $a_{2}$, $p_{2}$, and $s_{2}$ chosen such that the total energy is conserved, with $0<a_{2}<1$.
@@
A representative sample of the near- and far-from-equilibrium distribution functions and their corresponding collision kernels in the training data set is displayed in Fig. 2. Note that, instead of using the distribution function and its corresponding collision kernel as input and output, we choose $p^{3}(f-f_{\mathrm{eq}})$ as the former and $p^{3}C$ as the latter. We perform the multiplication by $p^{3}$ because we have observed improved energy conservation. Similarly, for the input data, we subtract the equilibrium distribution to facilitate the convergence towards the thermal fixed point.
@@ III.2 Hyperparameter optimization and uncertainty quantification
Many different neural network architectures have been described and proposed in the literature, which are designed for different tasks and purposes. Here, we use a feed-forward neural network with linear layers and ReLU functions as activation functions to obtain the collision kernel [^1] More precisely, if the output is modeled with a network of $N$ linear layers, it is obtained as $\mathrm{\mathbf{y}}=L_{N}(L_{N-1}(\cdots(L_{1}(\mathrm{\mathbf{x}}))))$. Here, $\mathrm{\mathbf{x}}$ is the input array and $L_{i}$ the $i$-th layer of the network. Therefore, the output of the $M$-th layer is computed by taking the output of the previous layer and applying the activation function: $y^{M}_{i}=\sigma\left(\sum_{j}W_{ij}y^{M-1}_{j}+b_{i}\right)$. In our case, we use a ReLU function as activation, which is defined as $\sigma(x)=\max(0,x)$.. For the precise architecture, i.e., number of layers and nodes per layer, we use an automated optimization framework, Ray Tune Liaw et al. (2018), to find the optimal set of hyperparameters. This hyperparameter tuning has the freedom to explore models with up to 5 layers; each of these can have 16, 32, 64, 128, 192, 256, 384, or 512 neurons.
We allow Ray Tune to choose not only the architecture of our neural network but also the learning rate [^2] The learning rate is the parameter used to quantify how much the gradient descent algorithm modifies the free parameters of the artificial neural network in each step of its training. Choosing an appropriate value is crucial to obtain a successful fit. as an additional hyperparameter. The architectures and learning rates with the best performance for each collision kernel are presented in Table 1. This tuning is performed for both collision kernels separately.
@@
Table 1: Best hyperparameters found for each collision kernels.
@@ III.3 Results
The training of the 20 neural networks, 10 for each collision kernel, which we performed for this case involved around 200000 samples, which translates to a volume of data of $\sim 100~\mathrm{MB}$. [^4] We provide the data in Ref. Barrera Cabodevila et al. (2025), which includes the raw data and preprocessing steps, effectively amounting to more than the $100~\mathrm{MB}$ of data. To evaluate the performance of the described approach, we compare the evolution of several initial conditions with the results obtained within the Monte Carlo approach. In Fig. 4 we present the distribution functions and their corresponding collision kernels at different times for three different evolutions. The first two are not contained in the training data, while the third one has been extracted from the training dataset. We observe excellent agreement between the two approaches. All the Monte Carlo results lie within the error bands obtained from the neural network evolution and Eq. (20).
@@ IV.4 Results
The first index $n\geq 0$ labels the power of $p$ in the integrand. The indices $l$ and $m$ are the indices belonging to the spherical harmonics $Y_{l}^{m}(\theta,\phi)$, and obey the condition $l\geq 0$ and $l\leq m\leq l$. Note that the Debye mass is proportional to the moment $M_{000}$, the number density to $M_{100}$, the energy density to $M_{200}$. Other familiar quantities can also be obtained from these moments. For instance, the pressure $P_{z}$ is given by $P_{z}/T^{4}=\frac{1}{3}M_{200}+\frac{2}{3}M_{220}$.
@@
We provide additional comparisons for simulations with different initial conditions in Fig. 7. There, we again first note the good overall agreement between the neural network and the Monte Carlo evolution. However, especially in the right panel, the error bars significantly grow during the evolution. Especially close to equilibrium, the network seems to struggle to remain and approach the thermal fixed point. Thus, despite enlarging the training data with a large amount of static distributions close to equilibrium, the neural network fails to accurately learn the approach to thermal equilibrium. This indicates that even more training data might be needed, or a different network architecture that by construction conserves the thermal fixed point. A promising possibility for enlarging the training dataset would be taking the predictions of the neural network close to equilibrium, and evolve them with the Monte Carlo method. This would generate more training data which complements the first dataset close to the thermal fixed point.
@@ Acknowledgements.
We would like to thank Andreas Ipp, David I. Müller, and Kayran Schmidt for insightful discussions. FL is supported by the Austrian Science Fund (FWF) under Grant DOI 10.55776/P34455. FL is a recipient of a DOC Fellowship of the Austrian Academy of Sciences at TU Wien (project 27203). The results in this paper have been achieved in part using the Austrian Scientific Computing (ASC) infrastructure, project 71444. SBC is supported by the European Research Council project ERC-2018-ADG-835105 YoctoLHC; by María de Maeztu grant CEX2023-001318-M and by project PID2023-152762NB-I00, both funded by MCIN/AEI/10.13039/-501100011033; from the Xunta de Galicia (CIGUS Network of Research Centres) and the European Union. SBC also acknowledges the support of Axudas de apoio á etapa predoutoral program (Ref. ED481A 2022/279).